3) Suppose you throw a baseball straight up at a velocity of 64 feet per second. A function can be created by expressing distance above the ground, s, as a function of time, t. This function is s = -16t2 + v0t + s0
• 16 represents 1/2g, the gravitational pull due to gravity (measured in feet per second2).
• v0 is the initial velocity (how hard do you throw the object, measured in feet per second).
• s0 is the initial distance above ground (in feet). If you are standing on the ground, then s0 = 0.
a) What is the function that describes this problem?
b) The ball will be how high above the ground after 1 second?
c) How long will it take to hit the ground?
d) What is the maximum height of the ball? What time will the maximum height be attained?

Answer by stanbon(75887)

a)s(t)=-16t^2+64t+0
b)s(1)=-16(1)^2+64(1)=-16+64=48 ft.
c)When it hits the ground the seight above the ground is zero.
-16t^2+64t=0
-16t(t-4)=0
t=0 or t=4
The ball is on the ground at time=0 and at time t=4
Cheers,
Stan H.